# Linguistic Collapse: Neural Collapse in (Large) Language Models

### v1

@@ Appendix G Interference with Scale – (Related to $\mathcal{NC}$2)
Appendix H Equiangularity with Scale – (Against $\mathcal{NC}$2)
@@
Figure 14: Variation in interference increases when scaling depth ($L$) in models trained for 1 (left) through 10 (right) with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom). Note this trend is against equiangularity, affirming the traditional $\mathcal{NC}$2 to be less useful than $\mathcal{GNC}$2 [16].
Appendix I Logarithmic Distances with Scale – (Related to $\mathcal{GNC}$2)
@@
Appendix J Hyperspherical Uniformity with Scale – $\mathcal{GNC}$2
@@
Appendix K Self-Duality with Scale – (Against $\mathcal{NC}$3)
Self-duality ($\mathcal{NC}3$) was originally the convergence of classifiers to class means after normalization:
@@
Instead, we use class-wise similarity (Equation 8) and its variation ($\mathcal{UNC}$3).
@@
Appendix L Uniformity Duality with Scale – $\mathcal{UNC}$3
@@
Appendix M Classifier Agreement – $\mathcal{NC}$4
@@
Appendix N Correlations with Generalization Performance
Figure 25: Validation loss shows some correlation with average mean norms (left) and their variations ($\mathcal{NC}_{2}$) (right).
Figure 26: Validation loss shows weak correlation with average interference and almost none with its variation (i.e. equiangularity, $\mathcal{NC}_{2}$) (right).
Figure 27: Validation loss shows some correlation with average kernel distances and with its variation (i.e. hyperspherical uniformity, $\mathcal{GNC}_{2}$) (right).

### v2

@@ Appendix H Interference with Scale – (Related to $\mathcal{NC}_{2}$)
Appendix I Equiangularity with Scale – (Against $\mathcal{NC}_{2}$)
@@
Figure 15: Variation in interference increases when scaling depth ($L$) in models trained for 1 (left) through 10 (right) with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom). Note this trend is against equiangularity, affirming the traditional $\mathcal{NC}_{2}$ to be less useful than $\mathcal{GNC}_{2}$ [19].
Appendix J Logarithmic Distances with Scale – (Related to $\mathcal{GNC}_{2}$)
@@
Appendix K Hyperspherical Uniformity with Scale – $\mathcal{GNC}_{2}$
@@
Appendix L Correlations of $\mathcal{(G)NC}_{2}$ with Generalization Performance
Figure 20: Generalization (validation loss) shows some correlation with average mean norms (left) and almost none with their variations (i.e. equinormness, $\mathcal{NC}_{2}$) (right).
Figure 21: Generalization (validation loss) correlated with average interference (left) and its variation (i.e. equiangularity, $\mathcal{NC}_{2}$) (centre). We also computed the empirical measure from [5] (right).
Figure 22: Validation loss shows some correlation with average (logarithmic) kernel distances and with their variation (i.e. hyperspherical uniformity, $\mathcal{GNC}_{2}$) (right).
Appendix M Self-Duality with Scale – (Against $\mathcal{NC}_{3}$)
Self-duality ($\mathcal{NC}_{3}$) was originally the convergence of classifiers to means up to rescaling [1]:
@@
Instead, we use class-wise cosine similarity (Equation 8) and its variation ($\mathcal{UNC}$3).
@@
Appendix N Uniformity Duality with Scale – $\mathcal{UNC}_{3}$
@@
Appendix O Correlations of $\mathcal{(U)NC}_{3}$ with Generalization Performance
Figure 27: Generalization (validation loss) correlated with average dot-product similarity (for interpretability purposes only) (left) and cosine similarity (classifier alignment, $\mathcal{NC}_{3}$) (centre). We also computed the empirical measure from [5] (right).
Figure 28: Generalization (validation loss) correlated with variation in dot-product similarity (for interpretability purposes only) (left) and cosine similarity (uniform duality, $\mathcal{UNC}_{3}$) (right).
Appendix P Classifier Agreement – $\mathcal{NC}_{4}$
@@
Appendix Q Examples for Interpretability
This section presents token-wise interpretability results found in top-layer embeddings from our most learned model (https://huggingface.co/rhubarbwu/TinyStories-12x1024_10L).
Table 4: Under TinyStories-12x1024_10L, these fifteen homonyms have much shorter mean embedding norms (i.e. closer to the global centre) than the average token. This is expected since homonyms typically present conflicts and interference.
GPT-Neo Token ID Token Text Scaled Norm (Eq. 5)
$808$ “row” $102.9727$
$1806$ “ring” $63.4789$
$2971$ “light” $74.4070$
$4053$ “well” $55.3457$
$4475$ “date” $111.7086$
$8664$ “bat” $83.0259$
$9464$ “left” $81.7467$
$15699$ “match” $114.7236$
$16469$ “spring” $80.0322$
$17796$ “bank” $90.8248$
$19204$ “wave” $43.5028$
$19836$ “close” $60.6004$
$22043$ “fair” $57.2620$
$28230$ “lead” $102.0583$
$36859$ “bowl” $88.7034$
(avg) (avg) $106.8762$
Table 5: Under TinyStories-12x1024_10L, the variability and interference of some English first names were far below those of the average token. This might be because names are distinct from one another and are not typically used in the same contexts as other words (aside from articles). The only names to have CDNV close to that of the average token are “Anna” and “Tim”. Note that the positive interference of the last row (average token) is not a typo.
GPT-Neo Token ID Token Text CDNV (Eq. 4) Interference (Eq. 6)
7371 Donald 0.0000817614 -0.008308
7554 John $00$0.007270
11006 David 0.0000924941 -0.006167
12041 Paul 0.0000966165 -0.006768
13256 Michael 0.0000885749 -0.006625
14731 James $00$0.006610
14967 Tim $00$0.004741
17121 William 0.0000933207 -0.005867
18050 Jim $00$0.006896
18308 Harry $00$0.006749
19156 Robert 0.0000813332 -0.006438
19206 Steve 0.0000962210 -0.006541
19962 Daniel $00$0.006963
20191 George $00$0.006674
20508 Andrew 0.0000928071 -0.006036
21868 Ryan 0.0000833253 -0.006594
22405 Thomas 0.0000952689 -0.006400
23865 Kevin 0.0000939218 -0.005817
24119 Mary 0.0000884645 -0.006322
24761 Brian 0.0000855408 -0.006003
24778 Martin 0.0000966354 -0.007183
25004 Eric 0.0000973125 -0.006565
25372 Matthew 0.0000810968 -0.005544
28711 Charles 0.0000787844 -0.006371
29284 Sarah 0.0000792671 -0.006366
30730 Luke $00$0.005830
31160 Anna $00$0.005005
32476 Henry $00$0.006685
32697 Anthony 0.0000659484 -0.006055
34831 Kelly 0.0000858015 -0.006017
40656 Robin $00$0.006808
42516 Kyle 0.0000815930 -0.005333
43187 Jennifer 0.0000875091 -0.006629
43568 Elizabeth 0.0000634704 -0.006462
43687 Laura 0.0000858422 -0.005922
44484 Alice 0.0000920828 -0.006416
45572 Jessica 0.0000840532 -0.005614
46751 Jacob 0.0000912490 -0.005349
AVERAGE AVERAGE 0.0001771000 0.000519